The Lenght of a rectangle is 8 inches shorter then 3 times its width. If the area of the rectangel is 35 square inches, find the dimensions of the rectangle.
Answer by sofiyac(983)

let width be x
then length is 3x-8
x(3x-8)=35
3x^2-8x-35=0
(3x+7)(x-5)=0
3x+7=0
3x=-7
x=-7/3 this won't work since we can't have negative length
x-5=0
x=5 so width is 5
then length is 3*5-8=7